# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ IV-A Stream 00 queuing system
The schematic for the Stream 00 queuing system is shown in Fig. 14. A Stream 00 queuing system $\overline{Z}$ has nodal arrival $a_{k}$ and non-nodal arrival $a_{\mathrm{O},k}$. At the source node, $a_{k}$ arrivals signify UAS being deployed on receiving requests at rate $\lambda$. Thus, $a_{k}$ is bernoulli distributed with arrival rate $\lambda$ and PGF $A(z)=(1-\lambda)+\lambda z$, whereas arrivals $a_{\mathrm{O},k}=0$ with probability 1 and PGF $A_{\mathrm{O}}(z)=1$. At any other node, the arrival distributions need to be computed considering departures from respective neighboring queuing systems. The Stream 00 queuing system has two $\beta_{0}$ queues. According to Rule 7, if a UAS is queuing, then it is in either of the two queues with probability $\eta$ and $1-\eta$. We merge the two queues and treat them as one. Similarly, following Rule 8, we merge the two $\widehat{\delta}_{\mathrm{O}}$ servers of Stream 00 queuing systems.
Figure 14: Stream 00 queuing system
@@
Let $v_{j,k}\in\{0,1,...,j\},\ j\in\{1,...,L\}$ be random variables such that at time slot $k$, $v_{j,k}$ gives the total number of UAS currently in the queue that have queued for no more than $j$ slots since entering the queue. $v_{1,k}=1$ implies that an arrival found the system busy in time slot $k-1$ and thus entered the queue and has queued for $1$ timeslot since then. $v_{L,k}$ gives the total number of UAS present in the queue. A UAS that has queued for $L+1$ slots (that is, $L$ successive transitions on $\beta_{0}$) exits the $\overline{Z}$ system in the next time slot. Hence, there is a queuing deadline of $L$ time slots to enter service, beyond which the UAS is lost from the system. As the queue never goes unbounded, the limiting probability distributions as $k\longrightarrow\infty$ exist.
@@
The UAS preference $\psi_{Z}\in\mathbb{Z}_{[-L,0]}$ decreases with each queued time slot. A UAS that has queued for the least number of slots has the highest preference. Hence, by Rule 3, if the $\overline{Z}$ system is idle ($Z_{\mathrm{UP}}$ is uncongested), then UAS with the highest preference enters service. This is fundamentally the Last Come, First Serve (LCFS) discipline. The following dynamical equations summarize the above queuing behavior when the system is busy and idle.
If the system is busy ($\theta_{k}=0$) then
$v_{1,k+1}=a_{k}$ (14)
$v_{j+1,k+1}=v_{j,k}+a_{k},\ \ j\in\{1,...,L-1\}$ (15)
If the system is idle ($\theta_{k}=1$) then following LCFS
$v_{1,k+1}=(a_{k}-1)^{+}=0$ (16)
$v_{j+1,k+1}=(v_{j,k}+a_{k}-1)^{+},\ \ j\in\{1,...,L-1\}$ (17)
@@
where the notation (.$(.)^{+}=\max\{0,.\}$. Eqn. 17 states that when $a_{k}=0$, then one UAS from the queue enters service (the one which arrived last). Hence, the number of UAS that have queued at most $j$ slots ($j\in\{1,...,L\}$) reduces by one. When $a_{k}=1$, then the arrival enters service, and the number of UAS that have queued $j$ slots by time slot $k$ would have queued $j+1$ slots by timeslot $k+1$. The arrivals $a_{k}$ and $v_{j,k}$ are independent at timeslot $k$. And,
$E[z^{(v_{j,k}+a_{k}-1)^{+}}]=V_{j}(0)+\left(\cfrac{V_{j}(z)-V_{j}(0)}{z}\right)A(z)$ (18)
The PGF $V_{j}(z),\ j\in\{1,...,L\}$ can be derived using the conditional PGF of Eqns. 14-17 as,
$V_{j}(z)$ $=E[z^{v_{j,k+1}}|\theta_{k}=0]P[\theta_{k}=0]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +E[z^{v_{j,k+1}}|\theta_{k}=1]P[\theta_{k}=1]$
Hence for $j\in\{1,...,L-1\}$,
$V_{1}(z)=A(z)\Theta_{0}+(1-\Theta_{0})$ (19)
$V_{j+1}(z)=\ A(z)V_{j}(z)\Theta_{0}+E[z^{(v_{j,k}+a_{k}-1)^{+}}](1-\Theta_{0})$ (20)
@@
By recursively substituting $V_{j}(z)$ in $V_{j+1}(z)$, we get $V_{L}(z)$. In the timeslot $k$, a UAS will enter service if and only if the system is uncongested and, more importantly, there are UAS available to enter service, that is
@@
    there is a UAS nodal arrival (that is $a_{k}=1$), or
@@
    there are UAS in queue (that is $v_{L,k}>0$), or
@@
    there is a UAS non-nodal arrival (that is $a_{\mathrm{O},k}=1$) available to descend towards $Z_{\mathrm{UP}}$ under the condition that there are no UAS nodal arrival and no UAS in the queue (descend condition),
Thus, as $k\longrightarrow\infty$, we have
$\sigma_{Z}=A(0)V_{L}(0)$ (21)
$\pi=(1-A(0))+A(0)(1-V_{L}(0))+\sigma_{Z}(1-A_{\mathrm{O}}(0))$ (22)
@@
where $\sigma_{Z}$ is referred to as steady-state descend probability that a UAS at node $\widetilde{Z}_{\mathrm{ON}}$ would descend into $Z_{\mathrm{UP}}$ zone and $\pi$ is the probability that a UAS is available to enter service in a given time slot. The PGF of the descend probability distribution is denoted by $B(z)=(1-\sigma_{Z})+z\sigma_{Z}$.
Eqn. 23 is the feedback relation that captures congestion being responsible for UAS entering the Stream 0 queue. Substituting Eqn. 23 in Eqn. 13, we get a polynomial equation in $\Theta_{0}$. There is an indeterminate term ($/0$ 0) in Eqn. 20, which we take advantage of using Rouché’s theorem. Knowing the steady state exists and $|z|<1$, according to Rouché’s theorem [29, 28], the roots of the denominator of Eqn. 20 are also roots of the numerator, which leads to pole-zero cancellations and the polynomial in $\Theta_{0}$ having a unique zero in $[0,1]$. The polynomial equation is solved numerically for $\Theta_{0}$. Substituting $\Theta_{0}$ in $\pi,\sigma_{Z},V_{j}(z),j\in\{1,L\}$, we get respective probabilities.
Note that the congestion definition in Eqn. 13 is defined solely on $M$ UAS being in the service or not. However, it does not account for when these $M$ UAS have entered service. It would be misleading to assume no correlation between the time instances the UAS enters service, as this affects the congestion period duration. For example, consider $M=2$ and service time $S=11$. If a UAS enters service in timeslot 00 followed by another UAS entering service in timeslot $1$. As $2$ UAS are in service, the $Z_{\mathrm{UP}}$ is congested in time slot $1$, and it remains congested for the next $S-1=10$ timeslots. Instead, if the second UAS had entered service in timeslot $10$, then the $Z_{\mathrm{UP}}$ turns congested in time slot $10$, but it remains congested only for the next $S-10=1$ timeslot. We introduce the above correlation by modeling congestion as a Markov modulated on/off regular process (MMRP) [30].
@@
IV-A1 Markov-modulated regular process
An MMRP (shown in Fig. 15(a)) is a two-state Markov chain, where the state of the Markov chain depends only on the state of a two-state background process. Here, $\theta_{k}\in\{0,1\}$ is the Markov state, with $\Theta_{0}$ and $\Theta_{1}=1-\Theta_{0}$ being the probability that $Z_{\mathrm{UP}}$ is congested and uncongested, respectively. The number of UAS ($n_{k}$) present in service is the background process with two states $n_{k}\geq M$ and $n_{k}<M$. If $n_{k}\geq M$ then $\theta_{k}=0$ and if $n_{k}<M$ then $\theta_{k}=1$. Let $\Theta_{10}=\mathrm{P}[n_{k+1}\geq M|n_{k}<M]=\mathrm{P}[\theta_{k+1}=0|\theta_{k}=1]$ is the transition probability that $Z_{\mathrm{UP}}$ if uncongested in timeslot $k$ turns congested in timeslot $k+1$, and $\Theta_{01}=\mathrm{P}[n_{k+1}<M|n_{k}\geq M]=\mathrm{P}[\theta_{k+1}=1|\theta_{k}=0]$ is the transition probability that $Z_{\mathrm{UP}}$ if congested in timeslot $k$ turns uncongested in timeslot $k+1$, respectively. $\Theta_{00}=(1-\Theta_{01})$ is the transition probability that $Z_{\mathrm{UP}}$ remains congested and $\Theta_{11}=(1-\Theta_{10})$ is the transition probability that it remains uncongested. Let $\pi_{e}=(\pi+\lambda_{e}-\pi\lambda_{e})$ denote the steady-state probability that a UAS is available to enter service (including exogenous arrivals) in any given time slot.
In a time interval of $S-1$ timeslots, if $M$ UAS were available to enter service, then the $Z_{\mathrm{UP}}$ would be congested in the $(S-1)$th timeslot. The $Z_{\mathrm{UP}}$ would remain congested if and only if no UAS are departing in the next time slot, that is, there are no UAS that have received $S-1$ slots of service as shown in Fig. 15(b). This behavior is summarized in Eqn 24.
$\Theta_{00}=\cfrac{\mathrm{P}[\theta_{S}=0,\theta_{S-1}=0]}{\mathrm{P}[\theta_{S-1}=0]}$
=($=\cfrac{{S-2\choose M}\left[\pi_{e}^{M}(1-\pi_{e})^{S-M-2}\right](1-\pi_{e})}{{S-1\choose M}\left[\pi_{e}^{M}(1-\pi_{e})^{S-M-1}\right]}=\cfrac{S-M-1}{S-1}$ (24)
In the $S-1$ timeslot interval, if no more than $M-1$ UAS were available to enter service, then the $Z_{\mathrm{UP}}$ would be uncongested in the $(S-1)$th timeslot. Given uncongested, the $Z_{\mathrm{UP}}$ would turn congested in the next time slot if $M-1$ UAS are already in service, no UAS are departing, and at least one UAS is available to enter the service (shown in Fig. 15(b)). This behavior is summarized in Eqn 25.
$\Theta_{10}$ $=\cfrac{\mathrm{P}[\theta_{S}=0,\theta_{S-1}=1]}{\mathrm{P}[\theta_{S-1}=1]}$
$=\cfrac{(\pi_{e})\left[{S-2\choose M-1}\pi_{e}^{M-1}(1-\pi_{e})^{S-M-1}\right](1-\pi_{e})}{\sum_{n=0}^{M-1}{S-1\choose n}\left[\pi_{e}^{n}(1-\pi_{e})^{S-n-1}\right]}$ (25)
The Markov modulated congestion probability as derived in [30] is given by,
@@ IV-A1 Markov-modulated regular process
IV-A2 Markov modulated Bernoulli process (MMBP)
In a given timeslot, whether a UAS would enter service or not is correlated to the $Z_{\mathrm{UP}}$ being congested and UAS being available. If $\theta_{k}=0$, then $w_{1,k}=0$. However, when $\theta_{k}=1$, if UAS is available then $w_{1,k}=1$ with probability $\pi$ else $w_{1,k}=0$ with probability $(1-\pi)$. $w_{1,k}$ is an MMBP, which refers to the Markov chain that generates a sequence of zeros when the background process is in one state and generates a Bernoulli process when the background process is in the second state (as shown in Fig. 16). Here, $w_{1,k}\in\{1,0\}$ is the two-state Markov chain with $\theta_{k}\in\{0,1\}$ being the background process. The Markov modulated probability that a UAS would enter service as derived in [30] is given by
$\mathrm{P}[w_{1,k}=1]=1-W_{1}^{*}(0)$ $=[\ \Theta_{11}\Theta_{1}^{*}+\Theta_{01}\Theta_{0}^{*}\ ]\pi$ (27)

### v2

@@ 4.1 Stream$\{0\}$ queuing system
The schematic for the Stream$\{0\}$ queuing system associated with Stream$(0)$ zone is shown in Fig. 11. A Stream$\{0\}$ queuing system $\overline{Z}$ has nodal arrival $a_{k}$ and non-nodal arrival $a_{\mathrm{O},k}$. At the source node, $a_{k}$ arrivals signify UAS being deployed on receiving requests at rate $\lambda$. Thus, $a_{k}$ is Bernoulli distributed with arrival rate $\lambda$ and PGF $A(z)=(1-\lambda)+\lambda z$, whereas arrivals $a_{\mathrm{O},k}=0$ with probability 1 and PGF $A_{\mathrm{O}}(z)=1$. At any other node, the arrival distributions need to be computed considering departures from respective neighboring queuing systems. The Stream$\{0\}$ queuing system has two $\gamma_{0}$ queues. According to Rule 8, if a UAS is queuing, then it is in either of the two queues with probability $\eta$ and $1-\eta$. We merge the two $\gamma_{0}$ queues and treat them as one. Similarly, following Rule 9, we merge the two $\widehat{\delta}_{\mathrm{O}}$ servers of Stream$\{0\}$ queuing system.
Figure 11: Stream$\{0\}$ queuing system
@@
Let $v_{j,k}\in\{0,1,...,j\},\ j\in\{1,...,L\}$ be random variables such that at timeslot $k$, $v_{j,k}$ gives the total number of UAS currently in the queue that have queued for no more than $j$ timeslots since entering the queue. Here, $v_{1,k}=1$ implies that an arrival found the system busy in timeslot $k-1$, so the arrival entered the queue and has queued for $1$ timeslot since then. $v_{L,k}$ gives the total number of UAS present in the queue. A UAS that has queued for $L$ timeslots (that is, executed $L$ successive transitions on $\gamma_{0}$ path) exits the $\overline{Z}$ system in the next timeslot. Hence, there is a queuing deadline of $L$ timeslots to enter service, beyond which the UAS is lost from the system. As the queue never goes unbounded, the limiting probability distributions as $k\longrightarrow\infty$ exist.
@@
The UAS preference $\psi_{Z}\in\mathbb{Z}_{[-L,0]}$ decreases with each queued timeslot. A UAS that has queued for the least number of timeslots has the highest preference. Hence, by priority (Rule 3), if the $\overline{Z}$ system is idle ($Z_{\mathrm{UP}}$ is uncongested), then UAS with the highest preference enters service. This is fundamentally the Last Come, First Serve (LCFS) discipline. The following dynamical equations summarize the queuing behavior when the system is busy and idle.
If the system is busy ($\theta_{k}=0$) then,
$v_{1,k+1}=a_{k}\ \ \ \ ,\ \ \ \ v_{j+1,k+1}=v_{j,k}+a_{k}\ \ \ \ ,\ \ \ \ j\in\{1,...,L-1\}$ (12)
If the system is idle ($\theta_{k}=1$) then following LCFS, we have
$v_{1,k+1}=(a_{k}-1)^{+}=0\ \ \ \ ,\ \ \ \ v_{j+1,k+1}=(v_{j,k}+a_{k}-1)^{+}\ \ \ \ ,\ \ \ \ j\in\{1,...,L-1\}$ (13)
@@
where the notation (.$(.)^{+}=\max\{0,.\}$. Eqn. 13 states that when $a_{k}=0$, then one UAS from the queue enters service (the one which arrived last). Hence, the number of UAS that have queued at most $j$ slots ($j\in\{1,...,L\}$) reduces by one. When $a_{k}=1$, then the arrival enters service, and the number of UAS that have queued $j$ slots by timeslot $k$ would have queued $j+1$ slots by timeslot $k+1$. The PGF $V_{j}(z),\ j\in\{1,...,L\}$ can be derived using the conditional PGF of Eqns. 12-13 as,
$V_{j}(z)\ =\ E[z^{v_{j,k+1}}|\theta_{k}=0]P[\theta_{k}=0]+E[z^{v_{j,k+1}}|\theta_{k}=1]P[\theta_{k}=1]$ (14)
The arrivals $a_{k}$ and $v_{j,k}$ are independent at timeslot $k$. Substituting the respective PGFs for $j\in\{1,...,L-1\}$, we get
$V_{1}(z)=A(z)\Theta_{0}+(1-\Theta_{0})\ \ \ \ ,\ \ \ \ V_{j+1}(z)=\ A(z)V_{j}(z)\Theta_{0}+\left[V_{j}(0)+\left(\cfrac{V_{j}(z)-V_{j}(0)}{z}\right)A(z)\right](1-\Theta_{0})$ (15)
@@
By recursively substituting $V_{j}(z)$ in $V_{j+1}(z)$, we get $V_{L}(z)$. In the timeslot $k$, a UAS will enter service if the system is uncongested and, more importantly, there are UAS available to enter service, that is,
@@
    there is a UAS nodal arrival (that is, $a_{k}=1$ with probability $(1-A(0))$), or
@@
    there are UAS in queue (that is, $v_{L,k}>0$ with probability $(1-V_{L}(0))$), or
@@
    there is a UAS non-nodal arrival (that is, $a_{\mathrm{O},k}=1$ with probability $(1-A_{\mathrm{O}}(0))$) available to descend towards $Z_{\mathrm{UP}}$ under the condition that there are no UAS nodal arrival and no UAS in the queue (descend condition),
As $k\longrightarrow\infty$, the probability $\pi$ that a UAS is available to enter service in a given timeslot is given by,
$\sigma_{Z}=A(0)V_{L}(0)\ \ \ ,\ \ \ \pi\ =\ (1-A(0))+A(0)(1-V_{L}(0))+\sigma_{Z}(1-A_{\mathrm{O}}(0))=\ (1-\sigma_{Z})+\sigma_{Z}(1-A_{\mathrm{O}}(0))$ (16)
where $\sigma_{Z}$ is referred to as the steady-state descend probability that a UAS at node $\widetilde{Z}_{\mathrm{ON}}$ would descend into $Z_{\mathrm{UP}}$ zone. The PGF of the descend probability distribution is denoted by $B(z)=(1-\sigma_{Z})+z\sigma_{Z}$. The steady-state probability that a UAS would enter service is,
@@
Eqn. 17 is the feedback relation that captures congestion being responsible for UAS entering the Stream$\{0\}$ queue. Substituting Eqn. 17 in Eqn. 11, we get a polynomial equation in $\Theta_{0}$ of degree $(L+1)(S-1)$. There is an indeterminate term ($/0$ 0) in Eqn. 15, which we take advantage of using Rouché’s theorem. Knowing the steady state exists and $|z|<1$, according to Rouché’s theorem (Kobayashi and Konheim (1977); Bruneel and Wuyts (1994)), the roots of the denominator of Eqn. 15 are also roots of the numerator, which leads to pole-zero cancellations and the polynomial in $\Theta_{0}$ having a zero in $[0,1]$ (existence proof provided in Theorem 1 in Appendix). The polynomial equation is solved numerically for $\Theta_{0}$. Substituting $\Theta_{0}$ in $\pi,\sigma_{Z},V_{j}(z),j\in\{1,L\}$, we get the respective probabilities.
Note that the congestion definition in Eqn. 11 is defined solely on $M$ UAS being in the service or not. However, it does not account for when these $M$ UAS have entered service. It would be misleading to assume no correlation between the time instances the UAS enters service, as this affects the congestion period duration. For example, assume $M=2$ and service time $S=11$ timeslots. Consider in timeslot ‘00’, $Z_{\mathrm{UP}}$ is uncongested, and a UAS enters service. If a second UAS enters service in timeslot ‘$1$’, then as $2$ UAS are in service, the $Z_{\mathrm{UP}}$ becomes congested in timeslot ‘$1$’. The first UAS will depart the system after receiving $11$ timeslots of service, and with the UAS departing, the $Z_{\mathrm{UP}}$ becomes uncongested in timeslot ‘$11$’. Thus, in a time interval of $11$ timeslots, the $Z_{\mathrm{UP}}$ is uncongested for one timeslot, and then it remains congested for the next $10$ timeslots. Instead, if the second UAS had entered service in timeslot ‘$10$’, then the $Z_{\mathrm{UP}}$ turns congested in timeslot ‘$10$’. By timeslot ‘$10$’, the first UAS would have already received $10$ timeslots of service and, in the next timeslot, would depart, leaving the system behind uncongested. In this case, the $Z_{\mathrm{UP}}$ is uncongested for $10$ timeslots, and it remains congested for only one timeslot. We introduce the above correlation by modeling congestion as a Markov modulated on/off regular process (MMRP) (Smiesko et al. (2023)).
@@
4.1.1 Markov-modulated regular process
An MMRP (shown in Fig. 12(a)) is a two-state Markov chain, where the state of the Markov chain depends only on the state of a two-state background process. Here, $\theta_{k}\in\{0,1\}$ is the Markov state, with $\Theta_{0}$ and $\Theta_{1}=1-\Theta_{0}$ being the probability that $Z_{\mathrm{UP}}$ is congested and uncongested, respectively. The number of UAS ($n_{k}$) present in service is the background process with two states $n_{k}\geq M$ and $n_{k}<M$. If $n_{k}\geq M$ then $\theta_{k}=0$ and if $n_{k}<M$ then $\theta_{k}=1$. Let $\Theta_{10}=\mathrm{P}[n_{k+1}\geq M|n_{k}<M]=\mathrm{P}[\theta_{k+1}=0|\theta_{k}=1]$ is the transition probability that $Z_{\mathrm{UP}}$, if uncongested in timeslot $k$ becomes congested in timeslot $k+1$, and $\Theta_{01}=\mathrm{P}[n_{k+1}<M|n_{k}\geq M]=\mathrm{P}[\theta_{k+1}=1|\theta_{k}=0]$ is the transition probability that $Z_{\mathrm{UP}}$, if congested in timeslot $k$ becomes uncongested in timeslot $k+1$, respectively. Here, $\Theta_{00}=(1-\Theta_{01})$ is the transition probability that $Z_{\mathrm{UP}}$ remains congested and $\Theta_{11}=(1-\Theta_{10})$ is the transition probability that it remains uncongested. Let $\pi_{e}=(\pi+\lambda_{e}-\pi\lambda_{e})$ denote the steady-state probability that a UAS is available to enter service (including exogenous arrivals) in any given timeslot.
If $M$ UAS were available to enter service in a time interval of $S-1$ timeslots, then the $Z_{\mathrm{UP}}$ would be congested in the $(S-1)$th timeslot. The $Z_{\mathrm{UP}}$ would remain congested if no UAS are departing in the next timeslot; that is, there are no UAS that have received $S-1$ slots of service as shown in Fig. 12(b). This event can be summarized in Eqn 18.
$\Theta_{00}\ =\ \cfrac{\mathrm{P}[\theta_{S}=0,\theta_{S-1}=0]}{\mathrm{P}[\theta_{S-1}=0]}\ =\ \cfrac{\left[{S-2\choose M}\pi_{e}^{M}(1-\pi_{e})^{S-M-2}\right](1-\pi_{e})}{\left[{S-1\choose M}\pi_{e}^{M}(1-\pi_{e})^{S-M-1}\right]}\ =\ \cfrac{S-M-1}{S-1}$ (18)
If $M-1$ UAS were available to enter service in the time interval of $S-1$ timeslots, then the $Z_{\mathrm{UP}}$ would be uncongested in the $(S-1)$th timeslot. Given $Z_{\mathrm{UP}}$ is uncongested, and if none of the $M-1$ UAS already in service departs, and at least one UAS is available to enter the service, then the $Z_{\mathrm{UP}}$ would become congested in the next timeslot (shown in Fig. 12(b)). This situation can be summarized in Eqn 19.
$\Theta_{10}\ =\ \cfrac{\mathrm{P}[\theta_{S}=0,\theta_{S-1}=1]}{\mathrm{P}[\theta_{S-1}=1]}\ =\ \cfrac{\pi_{e}\left[{S-2\choose M-1}\pi_{e}^{M-1}(1-\pi_{e})^{S-M-1}\right](1-\pi_{e})}{\sum_{n=0}^{M-1}\left[{S-1\choose n}\pi_{e}^{n}(1-\pi_{e})^{S-n-1}\right]}$ (19)
The Markov modulated congestion probability following MMRP derivations in Smiesko et al. (2023) is given by,
@@ 4.1.1 Markov-modulated regular process
4.1.2 Markov modulated Bernoulli process (MMBP)
In a given timeslot, whether a UAS would enter service or not is correlated to the $Z_{\mathrm{UP}}$ being congested and UAS being available. If $\theta_{k}=0$, then $w_{1,k}=0$. However, when $\theta_{k}=1$, if UAS is available then $w_{1,k}=1$ with probability $\pi$ else $w_{1,k}=0$ with probability $(1-\pi)$. Note that $w_{1,k}$ is an MMBP, which refers to the Markov chain that generates a sequence of zeros when the background process is in one state and generates a Bernoulli process when the background process is in the second state (as shown in Fig. 13). Here, $w_{1,k}\in\{1,0\}$ is the two-state Markov chain with $\theta_{k}\in\{0,1\}$ being the background process. The Markov modulated probability that a UAS would enter service following MMBP derivations in Smiesko et al. (2023) is given by
[…]